# Robustness of Dark Energy Phenomenology Across Different Parameterizations

### v1

@@ Abstract
The recent evidence for dynamical dark energy from DESI, in combination with other cosmological data, has generated significant interest in understanding the nature of dark energy and its underlying microphysics. However, interpreting these results critically depends on how dark energy is parameterized. This paper examines the robustness of conclusions about the viability of particular kinds of dynamical dark energy models to the choice of parameterization, focusing on four popular two-parameter families: the Chevallier-Polarski-Linder (CPL), Jassal-Bagla-Padmanabhan (JBP), Barboza-Alcaniz (BA), and exponential (EXP) parameterizations. We find that conclusions regarding the viability of minimally and non-minimally coupled quintessence models are independent of the parameterization adopted. We demonstrate this both by mapping these dark energy models into the $(w_{0},w_{a})$ parameter space defined by these various parameterizations and by showing that all of these parameterizations can equivalently account for the phenomenology predicted by these dark energy models to a high degree of accuracy.
@@ 1 Introduction
Understanding the nature of dark energy, alongside inflation and dark matter, is one of the three great mysteries of modern cosmology [1, 2, 3, 4, 5]. Recent results from the DESI collaboration, in combination with data from supernovae and CMB experiments, have provided the first statistically significant evidence that dark energy may be dynamical [6, 7, 8, 9, 10, 11]. Understandably, this has kicked off a wave of recent literature exploring these results and their implication for dark energy [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70].
Much of this literature has focused on how significantly our interpretation of the data is influenced or determined by what are, to some extent, arbitrary choices in terms of how we parameterize the dark energy equation of state. For example, the commonly used Chevallier-Polarski-Linder (CPL) parameterization, $w(a)=w_{0}+w_{a}\left(1-a\right)$, assumes that the dark energy equation of state evolves linearly with the scale factor and is given in terms of two parameters: the value of the equation of state now, $w_{0}$, and its time derivative, $w_{a}$ [71, 72]. Yet, recent work has pointed to certain issues or ambiguities that can arise when using it, such as its sensitivity to the redshifts probed by the data and the potential for misleading inferences when it is extrapolated outside the ranges of the data constraining it [73, 74, 75, 12]. Of course, there are no a priori reasons that guarantee that a linear form for the equation of state is sufficient to capture the relevant physics of dark energy and physicists have proposed other similar ways of parameterizing the dark energy equation of state in terms of two parameters with more intricate functional forms [76, 77, 78]. [^1] There are a tremendous variety of other options as well besides the most popular two parameter models considered here, including other two parameter models as well as one and three parameter models. See e.g. [79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92] and references therein for a small sample of this large literature. While recent studies have found that these too show a preference for dynamical dark energy [28, 93, 94], these proposals do not obviously resolve the ambiguities that have been pointed out in the case of the CPL parameterization. And finally, physicists have pursued more model agnostic methods, which have led to conflicting results depending on the approach adopted, with some favoring dynamical dark energy and others favoring $\Lambda$CDM (see e.g. [18, 29, 21, 17, 14]).
@@
This paper brings these strands of the literature together, focusing on the most popular two parameter families of dark energy parameterizations which include the CPL, Jassal-Bagla-Padmanabha (JBP), Barboza-Alcaniz (BA), and exponential (EXP) parameterizations. The questions we seek to answer are:
@@ 2 Parameterizing Dark Energy
$H^{2}(a)=H_{0}^{2}\left[\Omega_{r}a^{-4}+\Omega_{\mathrm{m}}a^{-3}+\Omega_{\mathrm{de}}e^{3\int_{a}^{1}\left(1+w_{\text{de }}\right)\mathrm{d}\ln\mathrm{a}}\right]$ (2.1)
@@ 3.1 Thawing quintessence with minimal coupling
where $\rho_{\varphi}$ and $P_{\varphi}$ are the scalar field energy density and pressure, respectively. The potential $V$ is the most significant factor in determining the evolution of $w(a)$ and the phenomenological behavior of dark energy. When $V$ dominates over the kinetic energy of the scalar field $\dot{\varphi}^{2}/2$, the equation of state approaches that of a cosmological constant $w(a)\simeq-1$. However, $w(a)$ will change over time due to the evolution of the scalar field $\varphi$. The dynamics of this evolution are governed by the scalar field’s equation of motion in a Friedmann background:
@@
$H^{2}\equiv\left(\frac{\dot{a}}{a}\right)^{2}=\frac{1}{3M^{2}_{pl}}\left[\frac{1}{2}\dot{\varphi}^{2}+V(\varphi)+\rho_{m}\right],$ (3.4)
with $\rho_{m}$ the energy density of the ordinary matter, respectively. The Hubble parameter today is $H_{0}=100h$ km s-1 Mpc-1, with $h\simeq 0.67$.
@@
This model can capture the entire phenomenology of thawing quintessence because, when we have a positive mass term in Eq. ((3.10)), it behaves as “slow-roll” thawing quintessence [109], which has been extensively studied and showed to converge to a universal phenomenological behavior given by an approximately linear evolution in the equation of state. This behavior is well-described by $dw/da|_{a=1}\approx 1.5[1+w(a=1)]$ (e.g. [109, 110, 101, 73, 111, 112]). When the mass is negative, this corresponds to “hilltop” quintessence [113, 114, 75, 73], which has much more varied dynamical behavior. If $\left|m^{2}\right|$ is small, the resulting $w(a)$ approximates the linear behavior of the positive branch. On the other hand, if $\left|m^{2}\right|$ is large, $w(a)\simeq-1$ for much of cosmic history, but can evolve very rapidly and non-linearly at more recent times depending on the choice of model parameters, allowing it to “sweep” over the $(w_{0},w_{a})$ parameter space. In addition to the above arguments, we can also understand why this model is able to subsume the phenomenology of thawing quintessence in the following way. As any arbitrary analytic potential can be expand in a Taylor series (e.g. [113, 73, 114]), the potential given in Eq. (3.10) can be seen as the leading order terms for a huge variety of scalar field potentials that broadly fall under either the “slow-roll” and “hilltop” types; and in the regime of field space for which quintessence can play the role of cosmological dark energy, these are the terms that will be relevant.
@@ 3.2 Thawing quintessence with non-minimal coupling
Another possibility for dark energy is to introduce a non-minimal coupling into the action Eq. (3.1), so that
@@
where $\xi$ is the dimensionless coupling coefficient between the scalar field $\varphi$ and gravity $R$. This introduces a modification into the gravitational sector (while keeping the matter sector of minimally coupled quintessence the same) [^2] Other kinds of non-minimal couplings such as a direct coupling between the scalar field and dark matter are also possible [115, 116, 117], but here we will restrict our focus only to non-minimal couplings with respect to gravity., which changes the corresponding scalar field equation of motion as well as the standard Friedmann equation from GR:
@@
$H^{2}=\frac{1}{3M_{\text{pl}}^{2}\left(1-\xi\varphi^{2}\right)}\left[\frac{1}{2}\dot{\varphi}^{2}+V(\varphi)+3H\xi\varphi\dot{\varphi}+\rho_{m}\right].$ (3.8)
@@
Interestingly, this modification can lead to a stable phantom crossing at recent times, which seems to be preferred by the expansion history data. As discussed in [13, 14, 46, 48, 55, 63, 64], dark energy models with a non-minimal coupling can improve the fit to the data over $\Lambda$CDM or quintessence. In particular, models that combine a non-minimal coupling like that in Eq. (3.6) with a “hilltop” potential can have a recent phantom crossing along with very sharp dynamical evolution, and they display a significant improvement to the fit of the data over $\Lambda$CDM or quintessence [13].
In this paper, to study these models we adopt a potential of the following form:
@@
where, as before, each of these terms can be thoughts of as the leading order terms in a Taylor expansion (but this time we keep the linear term for full generality). This model has been implemented in hi_class [107, 108], where the initial conditions are chosen such that $\varphi_{ini}=0$ and $\dot{\varphi}_{ini}\simeq 0$. We tune $V_{0}$ to satisfy the equations of motion, finding that $V_{0}\simeq 3H_{0}^{2}\Omega_{\varphi}$ is a robust starting point for the numerical solvers. Similar to before, the units of $V_{0}$ are given in $(H_{0}/h)^{2}$, the units of $\beta$ are given in $M_{\rm Pl}(H_{0}/h)^{2}$, and the units of $m^{2}$ are given in $M^{2}_{\rm Pl}(H_{0}/h)^{2}$. Of course, as a modified gravity model, this model can suffer from instabilities (see e.g. [118, 119]). In this paper we will be primarily concerned with predictions for background quantities, but we use hi_class to perform stability checks and ensure viability of all the parameter choices considered here by computing linear cosmological perturbations for each model.
In analogy with thawing quintessence above, we can refer to this as Non-minimal Thawing Quintessence as it will phenomenologically capture the behavior of dark energy models with this kind of non-minimal coupling and an arbitrary potential. For example, it captures the behavior of non-minimal models with an exponential potential and with a hilltop potential which have been studied in [13, 14, 46], as again this form can be seen as corresponding to a Taylor expansion in the potential. In [13], it was demonstrated that such models can saturate much of the thawing region of the CPL $(w_{0},w_{a})$ plane. While a full Bayesian analysis of this model is currently in preparation, here we will focus on the phenomenology and whether or not the aforementioned dark energy parametrizations are equipped to capture it.
@@ 3.3 Projecting dark energy models onto the $(w_{0},w_{a})$ plane
For non-minimal thawing quintessence, we generate $\simeq$ 10,000 models in the ranges $\xi\in[1.5,3.0]$, $\beta\in[0.0,10.0]$, and $m^{2}\in[-10.0,2.0]$. This flexibility, however, allows the model to cover a very wide range of behaviors for $w(a)$, including large parts of the phantom region today ($w_{0}<-1$) and huge swaths of the thawing region ($w_{a}<0$). However, for the purposes of this paper, the range of parameters was chosen to reproduce the slice of parameter space that most closely corresponds to the current constraints on the equation of state. In all cases considered here, we have validated that the number of models sampled is sufficient as the results do not change with the addition of more models.
@@ 4 Data and likelihood
We constrain the cosmological parameters and $({w_{0},w_{a}})$ using the Metropolis-Hasting algorithm [123, 124, 125, 126] and data likelihoods as implemented in Cobaya. We sample the cosmological parameters $\{\omega_{b},\Omega_{m},H_{0},n_{s},\ln 10^{10}A_{s},\tau\}$, together with $(w_{0},w_{a})$, with the priors shown in Table (1). We use the Boltzmann code hi_class to compute the observables and assess convergence with the Gelman-Rubin criterion [127], requiring $R-1<0.02$.
@@ 5.1 Dark energy models in the $(w_{0},w_{a})$ plane
Figure 1: Equation of state, $w(a)$, for a representative dark energy models from both thawing quintessence (left) and nonminimal thawing quintessence (right), alongside the corresponding best fit CPL, JBP, EXP, and BA models when these are determined by fitting directly to the compressed BAO, SNe, and CMB data using Eq. (3.11).
We start by showing how the different parametrizations represent the equation of state. Fig. (1) shows the equation of state for a thawing quintessence and non-minimal thawing quintessence model and the corresponding best fitting parametrizations. The dark energy models have cosmological parameters $\Omega_{m}=0.315$, $H_{0}=67.5$ and dark energy parameters $(V_{0},m^{2})=(0.975,-16.0)$ for thawing quintessence and $(\xi,V_{0},\beta,m^{2})=(1.5,0.83,2.0,-2.0)$, for the non-minimal case. The best-fitting parameters for the thawing quintessence model are CPL: $(w_{0},w_{a})=(-0.94,-0.12)$, JBP: $(w_{0},w_{a})=(-0.92,-0.34)$, EXP: $(w_{0},w_{a})=(-0.95,-0.09)$, and BA: $(w_{0},w_{a})=(-0.95,-0.06)$; and for the non-minimal model: CLP: $(w_{0},w_{a})=(-0.91,-0.30)$, JBP: $(w_{0},w_{a})=(-0.87,-0.73)$, EXP: $(w_{0},w_{a})=(-0.92,-0.23)$, and BA: $(w_{0},w_{a})=(-0.92,-0.15)$. While the various dark energy parameterizations all produce reasonable looking fits (again determined by fitting the observables directly) when plotted against the equation of state, they can also be misleading when extrapolated over all of cosmic history as the results will depend on the redshifts probed by the data and their associated uncertainties (see [73, 13, 75, 12] for more related discussion). That is, $w_{0}$ often differs substantially from $w(a=1)$ (even though it is often interpretted as being the value of the equation of state today) because the best fit $(w_{0},w_{a})$ is determined by fitting over a range of observations mostly from the distant past, meaning that its extrapolation to the present time might not be faithful to the actual model under consideration. Likewise, the parameter values suggest very different behavior at $w(a=0)$ (the very distant past) than is seen in the actual model for all parameterizations considered. Therefore, the best fitting parameters could suggest unphysical behavior. For example, extrapolating the CPL, EXP, or BA fits over all of cosmic history for the thawing quintessence model would lead one to conclude that the dark energy model it describes is phantom, which is clearly not the case as the actual model can never be phantom or cross the phantom line. In other words, it is important to not over-interpret any parameterization for dark energy as they can be misleading, even if they reproduce the observables of the dark energy models quite closely (more on this in Section (5.2)).

### v2

@@ Abstract
The recent evidence for dynamical dark energy from DESI, in combination with other cosmological data, has generated significant interest in understanding the nature of dark energy and its underlying microphysics. However, interpreting these results critically depends on how dark energy is parameterized. This paper examines the robustness of conclusions about the viability of particular kinds of dynamical dark energy models to the choice of parameterization, focusing on four popular two-parameter models: the Chevallier-Polarski-Linder (CPL), Jassal-Bagla-Padmanabhan (JBP), Barboza-Alcaniz (BA), and exponential (EXP) parameterizations. We find that conclusions regarding the viability of minimally and non-minimally coupled quintessence models are independent of the parameterization adopted. We demonstrate this both by mapping these dark energy models into the $(w_{0},w_{a})$ parameter space defined by these various parameterizations and by showing that all of these parameterizations can equivalently account for the phenomenology predicted by these dark energy models to a high degree of accuracy.
@@ 1 Introduction
Understanding the nature of dark energy, alongside inflation and dark matter, is one of the three great mysteries of modern cosmology [1, 2, 3, 4, 5]. Recent results from the DESI collaboration, in combination with data from supernovae and CMB experiments, have provided the first statistically significant evidence that dark energy may be dynamical [6, 7, 8, 9, 10, 11]. Understandably, this has kicked off a wave of recent literature exploring these results and their implication for dark energy [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79].
Much of this literature has focused on how significantly our interpretation of the data is influenced or determined by what are, to some extent, arbitrary choices in terms of how we parameterize the dark energy equation of state. For example, the commonly used Chevallier-Polarski-Linder (CPL) parameterization, $w(a)=w_{0}+w_{a}\left(1-a\right)$, assumes that the dark energy equation of state evolves linearly with the scale factor and is given in terms of two parameters: the value of the equation of state now, $w_{0}$, and its time derivative, $w_{a}$ [80, 81]. Yet, recent work has pointed to certain issues or ambiguities that can arise when using it, such as its sensitivity to the redshifts probed by the data and the potential for misleading inferences when it is extrapolated outside the ranges of the data constraining it [82, 83, 84, 12]. Of course, there are no a priori reasons that guarantee that a linear form for the equation of state is sufficient to capture the relevant physics of dark energy and physicists have proposed other similar ways of parameterizing the dark energy equation of state in terms of two parameters with more intricate functional forms [85, 86, 87]. [^1] There are a tremendous variety of other options as well besides the most popular two parameter models considered here, including other two parameter models as well as one and three parameter models. See e.g. [88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102] and references therein for a small sample of this large literature. While recent studies have found that these too show a preference for dynamical dark energy [29, 103, 104, 105], these proposals do not obviously resolve the ambiguities that have been pointed out in the case of the CPL parameterization. And finally, physicists have pursued more model agnostic methods, which have led to conflicting results depending on the approach adopted, with some favoring dynamical dark energy and others favoring $\Lambda$CDM (see e.g. [18, 30, 22, 17, 14]).
@@
This paper brings these strands of the literature together, focusing on the most popular two-parameter dark energy parameterizations which include the CPL, Jassal-Bagla-Padmanabha (JBP), Barboza-Alcaniz (BA), and exponential (EXP) parameterizations. The questions we seek to answer are:
@@ 2 Parameterizing Dark Energy
$H^{2}(a)=H_{0}^{2}\left[\Omega_{\mathrm{r}}a^{-4}+\Omega_{\mathrm{m}}a^{-3}+\Omega_{\mathrm{de}}e^{3\int_{a}^{1}\left(1+w_{\text{de }}\right)\mathrm{d}\ln\mathrm{a}}\right]$ (2.1)
@@ 3.1 Thawing quintessence with minimal coupling
where $\rho_{\varphi}$ and $P_{\varphi}$ are the scalar field energy density and pressure, respectively. The potential $V(\varphi)$ is the most significant factor in determining the evolution of $w(a)$ and the phenomenological behavior of dark energy. When $V(\varphi)$ dominates over the kinetic energy of the scalar field $\dot{\varphi}^{2}/2$, the equation of state approaches that of a cosmological constant $w(a)\simeq-1$. However, $w(a)$ will change over time due to the evolution of the scalar field $\varphi$. The dynamics of this evolution are governed by the scalar field’s equation of motion in a Friedmann background:
@@
$H^{2}\equiv\left(\frac{\dot{a}}{a}\right)^{2}=\frac{1}{3M^{2}_{\rm Pl}}\left[\frac{1}{2}\dot{\varphi}^{2}+V(\varphi)+\rho_{\rm m}\right],$ (3.4)
with $\rho_{\rm m}$ the energy density of the ordinary matter, respectively. The Hubble parameter today is $H_{0}=100h$ km s-1 Mpc-1, with $h\simeq 0.67$.
@@
This model can capture the entire phenomenology of thawing quintessence because, when we have a positive mass term in Eq. (3.11), it behaves as “slow-roll” thawing quintessence [120], which has been extensively studied and showed to converge to a universal phenomenological behavior given by an approximately linear evolution in the equation of state. This behavior is well-described by $dw/da|_{a=1}\approx 1.5[1+w(a=1)]$ (e.g. [120, 121, 112, 82, 122, 123]). When the mass is negative, this corresponds to “hilltop” quintessence [124, 125, 84, 82, 75], which has much more varied dynamical behavior. If $\left|m^{2}\right|$ is small, the resulting $w(a)$ approximates the linear behavior of the positive branch. On the other hand, if $\left|m^{2}\right|$ is large, $w(a)\simeq-1$ for much of cosmic history, but can evolve very rapidly and non-linearly at more recent times depending on the choice of model parameters, allowing it to “sweep” over the $(w_{0},w_{a})$ parameter space. In addition to the above arguments, we can also understand why this model is able to subsume the phenomenology of thawing quintessence in the following way. As any arbitrary analytic potential can be expand in a Taylor series (e.g. [124, 82, 125]), the potential given in Eq. (3.11) can be seen as the leading order terms for a huge variety of scalar field potentials that broadly fall under either the “slow-roll” and “hilltop” types; and in the regime of field space for which quintessence can play the role of cosmological dark energy, these are the terms that will be relevant.
@@ 3.2 Thawing quintessence with non-minimal coupling
Another possibility for dark energy is to introduce a non-minimal coupling into the action:
@@
where $\xi$ is the dimensionless coupling coefficient between the scalar field $\varphi$ and gravity $R$. This introduces a modification into the gravitational sector (while keeping the matter sector of minimally coupled quintessence the same) [^2] Other kinds of non-minimal couplings such as a direct coupling between the scalar field and dark matter are also possible (e.g. [126, 127, 128, 129, 76]), but here we will restrict our focus only to non-minimal couplings with respect to gravity., which changes the corresponding scalar field equation of motion as well as the standard Friedmann equation from GR:
@@
$H^{2}=\frac{1}{3M_{\text{\rm Pl}}^{2}\left(1-\xi\varphi^{2}\right)}\left[\frac{1}{2}\dot{\varphi}^{2}+V(\varphi)+3H\xi\varphi\dot{\varphi}+\rho_{m}\right].$ (3.8)
@@
Interestingly, this modification can lead to a stable phantom crossing at recent times, which seems to be preferred by the expansion history data. As discussed in [13, 14, 47, 49, 56, 64, 65], dark energy models with a non-minimal coupling can improve the fit to the data over $\Lambda$CDM or quintessence. In particular, models that combine a non-minimal coupling like that in Eq. (3.6) with a “hilltop” potential,
$V(\varphi)=V_{0}\left(1-\beta e^{-\lambda\varphi}\right),$ (3.10)
can have a recent phantom crossing along with very sharp dynamical evolution, and they display a significant improvement to the fit of the data over $\Lambda$CDM or quintessence [13].
In this paper, we instead adopt a more “agnostic” potential of the following form:
@@
where, as before, each of these terms can be thoughts of as the leading order terms in a Taylor expansion. This model has been implemented in hi_class [118, 119], where the initial conditions are chosen such that $\varphi_{ini}=0$ and $\dot{\varphi}_{ini}\simeq 0$. We tune $V_{0}$ to satisfy the equations of motion, finding that $V_{0}\simeq 3H_{0}^{2}\Omega_{\varphi}$ is a robust starting point for the numerical solvers. Similar to before, the units of $V_{0}$ are given in $M^{2}_{\rm Pl}(H_{0}/h)^{2}$, the units of $\beta$ are given in $M_{\rm Pl}(H_{0}/h)^{2}$, and the units of $m^{2}$ are given in $(H_{0}/h)^{2}$. Of course, as a modified gravity model, this model can suffer from instabilities (see e.g. [130, 131]). In this paper we will be primarily concerned with predictions for background quantities, but we use hi_class to perform stability checks and ensure viability of all the parameter choices considered here by computing linear cosmological perturbations for each model.
In analogy with thawing quintessence above, we can refer to this as Non-Minimal Thawing Quintessence as it will phenomenologically capture the behavior of dark energy models with this kind of non-minimal coupling and an arbitrary potential. For example, it captures the behavior of non-minimal models with an exponential potential and with a hilltop potential which have been studied in [13, 14, 47], as again this form can be seen as corresponding to a Taylor expansion in the potential. In [13], it was demonstrated that such models can saturate much of the thawing region of the CPL $(w_{0},w_{a})$ plane. See also [20] for a full Bayesian analysis where it is found that this type of non-minimally coupled dark energy model is strongly favoured over $\Lambda$CDM. However, in this paper, we will focus on the phenomenology and whether or not the aforementioned dark energy parametrizations are equipped to capture it.
@@ 3.3 Projecting dark energy models onto the $(w_{0},w_{a})$ plane
For non-minimal thawing quintessence, we generate $\simeq$ 10,000 models in the ranges $\xi\in[1.5,3.0]$, $\beta\in[0.0,10.0]$, and $m^{2}\in[-10.0,2.0]$. The flexibility of this model; however, allows it to cover a very wide range of behaviors for $w(a)$, including large parts of the phantom region today ($w_{0}<-1$) and huge swaths of the thawing region ($w_{a}<0$). However, for the purposes of this paper, the range of parameters was chosen to reproduce the slice of parameter space that most closely corresponds to the current constraints on the equation of state. In all cases considered here, we have validated that the number of models sampled is sufficient as the results do not change with the addition of more models.
@@ 4 Data and likelihood
We constrain the cosmological parameters and $({w_{0},w_{a}})$ using the Metropolis-Hasting algorithm [135, 136, 137, 138] and data likelihoods as implemented in Cobaya. We sample the cosmological parameters $\{\omega_{b},\Omega_{m},H_{0},n_{s},\ln 10^{10}A_{s},\tau\}$, together with $(w_{0},w_{a})$, with the priors shown in Table (1) (while keeping the neutrino mass fixed to $\sum m_{\nu}=0.06\,\mathrm{eV}$ throughout). We use the Boltzmann code hi_class to compute the observables and assess convergence with the Gelman-Rubin criterion [139], requiring $R-1<0.02$.
@@ 5.1 Dark energy models in the $(w_{0},w_{a})$ plane
Figure 1: Equation of state, $w(a)$, for representative dark energy models from both thawing quintessence (left) and non-minimal thawing quintessence (right), alongside the corresponding best fit CPL, JBP, EXP, and BA models when these are determined by fitting directly to the compressed BAO, SNe, and CMB data using Eq. (3.12).
We start by showing how the different parametrizations represent the equation of state. Fig. (1) shows the equation of state for a representative thawing quintessence and a representative non-minimal thawing quintessence model and the corresponding best fitting parametrizations. The dark energy models have cosmological parameters $\Omega_{m}=0.315$, $H_{0}=67.5$ and dark energy parameters $(V_{0},m^{2})=(0.975,-16.0)$ for thawing quintessence and $(\xi,V_{0},\beta,m^{2})=(1.5,0.83,2.0,-2.0)$, for the non-minimal case. The best-fitting parameters for the thawing quintessence model are: CPL: $(w_{0},w_{a})=(-0.94,-0.12)$, JBP: $(w_{0},w_{a})=(-0.92,-0.34)$, EXP: $(w_{0},w_{a})=(-0.95,-0.09)$, and BA: $(w_{0},w_{a})=(-0.95,-0.06)$; and for the non-minimal model: CLP: $(w_{0},w_{a})=(-0.91,-0.30)$, JBP: $(w_{0},w_{a})=(-0.87,-0.73)$, EXP: $(w_{0},w_{a})=(-0.92,-0.23)$, and BA: $(w_{0},w_{a})=(-0.92,-0.15)$. While the various dark energy parameterizations all produce reasonable looking fits (again determined by fitting the observables directly) when plotted against the equation of state, they can also be misleading when extrapolated over all of cosmic history as the results will depend on the redshifts probed by the data and their associated uncertainties (see [82, 13, 84, 12] for more related discussion). That is, $w_{0}$ often differs substantially from $w(a=1)$ (even though it is often interpretted as being the value of the equation of state today) because the best fit $(w_{0},w_{a})$ is determined by fitting over a range of observations mostly from the distant past, meaning that its extrapolation to the present time might not be faithful to the actual model under consideration. Likewise, the parameter values suggest very different behavior at $w(a=0)$ (the very distant past) than is seen in the actual model for all parameterizations considered. Therefore, the best fitting parameters could suggest unphysical behavior. For example, extrapolating the CPL, EXP, or BA fits over all of cosmic history for the thawing quintessence model would lead one to conclude that the dark energy model it describes is phantom, which is clearly not the case as the actual model can never be phantom or cross the phantom line. In other words, it is important to not over-interpret any parameterization for dark energy as they can be misleading, even if they reproduce the observables of the dark energy models quite closely (more on this in Section (5.2)).